### v1

# VoCo-LLaMA: Towards Vision Compression with Large Language Models

Abstract
Vision-Language Models (VLMs) have achieved remarkable success in various multi-modal tasks, but they are often bottlenecked by the limited context window and high computational cost of processing high-resolution image inputs and videos. Vision compression can alleviate this problem by reducing the vision token count. Previous approaches compress vision tokens with external modules and force LLMs to understand the compressed ones, leading to visual information loss. However, the LLMs’ understanding paradigm of vision tokens is not fully utilised in the compression learning process. We propose VoCo-LLaMA, the first approach to compress vision tokens using LLMs. By introducing Vision Compression tokens during the vision instruction tuning phase and leveraging attention distillation, our method distill how LLMs comprehend vision tokens into their processing of VoCo tokens. VoCo-LLaMA facilitates effective vision compression and improves the computational efficiency during the inference stage. Specifically, our method achieves minimal performance loss with a compression ratio of 576, resulting in up to 94.8% fewer FLOPs and 69.6% acceleration in inference time. Furthermore, through continuous training using time-series compressed token sequences of video frames, VoCo-LLaMA demonstrates the ability to understand temporal correlations, outperforming previous methods on popular video question-answering benchmarks. Our approach presents a promising way to unlock the full potential of VLMs’ contextual window, enabling more scalable multi-modal applications. The project page, along with the associated code, can be accessed via the following link.
1 Introduction
The advent of visual-language models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] has led to significant advancements in visual understanding. Particularly, high-resolution image encoding [7, 8] and the incorporation of more video frames [9, 10] have been shown to enhance the capabilities of both large visual-language models and large video-language models, respectively. However, the large number of vision tokens occupies a substantial portion of the valuable context window of the large language model, leading to expensive computational costs, as shown in Figure 1 (a). For instance, when using high-resolution image inputs in LLaVA-1.6 [7], a single image with a resolution of 672 672 is divided into four smaller patches, each encoded with a 336 336 resolution input. This process yields an image representation consisting of 2304 vision tokens, occupying over half of the context length. Furthermore, as the number of input images increases, the context window for text will be further bottlenecked. For example, Vicuna-1.5 [11] can only handle approximately 7 frames (7 576 = 4032 tokens) within its 4k context length, given away the text inputs. [9, 10] investigate the efficacy of extending the context length to the million-level mark to mitigate this issue, but this approach necessitates expensive computational resources (e.g., [9] requires over 1000 v4 TPUs) and engineering efforts in data preparation and framework development.
Effective representation of visual information within limited context length is crucial for addressing the challenges. A promising approach is to reduce the number of vision tokens representing an image, ideally to just a handful or even a single token. As illustrated in Figure 1 (b), previous methods [1, 3, 4, 6, 12, 13] have exploited the use of Q-Former [1] or straight-forward pooling operations to "compress" the encoded vision tokens. These methods compress the vision tokens with external modules and force LLMs to understand the compressed ones. This may disrupt the established paradigm of how LLMs comprehend original uncompressed vision tokens and bring loss of visual information. However, LLMs with uncompressed vision token input have already possessed visual understanding capabilities. Therefore, a more intuitive approach of vision compression is to distill how LLMs understand vision tokens into their processing of compressed tokens.
We propose VoCo-LLaMA, the first vision compression method that exploits the inherent functionality of large language models. As demonstrated in Figure 1 (c), we introduce Vision Compression (VoCo) tokens between visual and text tokens. By simply modifying the attention mechanism, we first isolate the visual and text tokens by inserting VoCo tokens. Subsequently, we establish an exclusive interaction pathway between the visual and text tokens via VoCo tokens. This facilitates the LLM itself to compress and distill the parsing vision tokens, specifically the transformer activations on top of them, into compact VoCo tokens. Benefiting from this compression strategy, our method achieves modality alignment without the specialized text-vision cross-modal fusion module design [1, 12].
Our method can learn vision compression through standard visual instruction tuning, eliminating the need for incurring additional training phase. Furthermore, when handling multiple repetitive visual inputs in the inference stage, VoCo-LLaMA enables caching and reusing compressed transformer activations, improving computational efficiency and reducing storage requirements for caching. VoCo-LLaMA can achieve a 576 compression rate (compressing 576 vision tokens into a single VoCo token) while maintaining 83.7% performance. Additionally, we observe a reduction in inference computation costs, with up to 99.8% in cache storage, 94.8% FLOPs and 69.6% in inference time.
To further investigate the efficacy of our proposed method in handling video input, we utilize the time-series VoCo token sequence obtained by compressing video frames to further explore the temporal modeling capabilities of VoCo-LLaMA. Without any additional designs, VoCo-LLaMA outperforms existing vision compression methods on common video question-answering benchmarks, with each video frame compressed into an equal number of tokens.
2 Related Work
LLMs and Text Compression. In recent years, large language models (LLMs) have sparked a technological revolution. As the scale of training data and model size continue to expand, models [11, 14, 15, 16, 17, 18, 19, 20] have demonstrated exceptional capabilities in understanding and generating language, showcasing the powerful abilities of LLMs in handling complex language tasks. In particular, models such as the LLaMA series [17, 18, 20], Vicuna [11], and Mistral [19] have emerged as foundational models or main components in many research works. However, the limited context window size in LLMs has long been a widely discussed topic in the natural language processing field. Compressing text to conserve valuable window length has been proven to be an efficient approach. Long-standing works, including [21, 22, 23, 24, 25], focus on storing text representations in transformers to achieve dense information representation. [26, 27] have demonstrated the effectiveness of distilling long text information into prompt-free student models. [28] has made significant progress in compressing information in long contexts by learning soft prompts. In a similar vein, recent studies [29, 30, 31] have explored the potential applications of compressing text in large language models. Although the context length limitation is a likewise pervasive challenge in vision-language models, the discussion of compressing visual information has been relatively understudied compared to the language model domain. Our work pioneers the use of LLMs’ learning capabilities to compress and distill information in vision tokens, aiming to bridging this gap in the field of VLMs.
VLMs and Vision Compression. The success of LLMs has inspired significant progress in vision language models (VLMs). By integrating visual encoders with LLMs, VLMs can effectively achieve cross-modal understanding and other visual language tasks through modality alignment and instruction tuning. Previous methods [1, 2, 3, 4, 5, 6, 7, 8, 13, 32] have substantiated the success of this training paradigm in visual understanding. The successful application of VLMs on images has also been rapidly extended to the video domain [9, 10, 12, 33, 34, 35, 36, 37, 38, 39]. With the input of higher-resolution images [7, 8] and more video frames [9, 10], VLMs can capture rich visual information for understanding and generation. However, as the number of vision tokens representing an input image increases, they take up a significant portion of the limited context window of language models, and can even exceed it. To address this, previous methods [1, 3, 4, 6, 13] have largely employed Q-Former [1], which maps images to fixed-length tokens in the language embedding space through learnable queries, compressing visual information. A more recent approach [12] has applied average pooling with a learnable linear layer to compress visual features through multi-stage training strategy. Although these methods perform moderately well at lower compression multiples, they cause a significant loss of valuable visual information when the number of compressed tokens reduces. VoCo-LLaMA distills the approach of LLMs in understanding vision tokens into their processing of compressed tokens, significantly reducing information loss during the vision compression process.
3 Method
We first introduce VoCo-LLaMA, a large language model capable of compressing lengthy vision tokens into compact VoCo tokens through attention distillation, which enables the efficient representation of visual information. Then, we build upon these compressed tokens to continue training VoCo-LLaMA, enabling our model to capture temporal dependencies within video data.
3.1 Vision Compression
Given a paired image and text input, we follow the design of most vision-language models (VLMs) and encode the image into a sequence of vision tokens , where is the number of the output patches from the visual encoder. Similarly, the text input is encoded into a sequence of text tokens . Consider an original, unmodified standard vision-language model (denoted as ), exemplified by LLaVA [2], depicted in Figure 2 (a). During visual instruction tuning, leverages both vision tokens and text tokens to predict the output through self-regression, and learns the distribution . For image compression models, our goal is to employ a compact set of compressed tokens to efficiently represent the sparse vision token set . Additionally, we aim to generate outputs that closely approximates the outputs of the original model when presented with identical image and text inputs.
Specifically speaking, with an image encoded as vision tokens , we formulate the image compression distillation process as learning a compression model that generates the output conditioned on the compressed tokens and the text tokens . This is achieved by learning a probability distribution . The optimization objective of is to minimize the loss function:
With above distillation objective, how to further distill the information within the vision tokens into the compressed token is the key of vision compression.
3.2 VoCo-LLaMA
As illustrated in Figure 2 (b), we propose VoCo-LLaMA, which introduces special Vision Compression (VoCo) tokens to leverage the LLM’s ability to compress and understand the compression representations of images. The input sequence to the large language model is formed by concatenating the vision tokens, the special tokens, and the text tokens, yielding a sequence:
In the training phase, we employ a two-stage attention mechanism. Initially, we impose a constraint on the text tokens, explicitly preventing them from attending to the original vision tokens, while forcing them to exclusively attend to the compressed and distilled VoCo tokens. Subsequently, the vision tokens are subjected to continuous attention from the VoCo tokens due to the casual transformer. This deliberate design ensures that the text tokens solely capture the distilled visual information encoded in the VoCo tokens, rather than directly interacting with the original vision tokens, achieving compressing distillation from vision tokens to compressed tokens.
The compression process of VoCo-LLaMA can be elegantly implemented by strategically modifying the attention mask. Specifically, we set the attention weights between the text tokens and the vision tokens to , effectively rendering the text tokens "isolated" to the vision tokens. Formally, let denote the attention mask, where if token attends to token , and otherwise. We define the modified attention mask as:
It is worth noting that our modification to the attention mask is based on the casual attention mask implementation in decoder-only models (e.g. LLaMA [17]), which is initialized as a lower-left triangular matrix. In practice, VoCo-LLaMA can be effectively trained under the standard supervised fine-tuning paradigm, leveraging the abundant image-text data readily available in VLMs. Furthermore, the VoCo token can be compactly represented as a set of transformer activations, allowing them to be cached to enhance inference efficiency.
VoCo-LLaMA enables the large language models to learn the compression process of vision tokens, , while simultaneously learning to understand the compressed VoCo tokens, i.e., learning the distribution . With VoCo-LLaMA, the optimization objective in Equation 2 can be further defined as:
Moreover, our design affords an additional advantage: the VoCo tokens exclusively interact with the input text during the compression process, establishing a seamless pathway between the visual and text modalities. Consequently, this enables us to effortlessly align the compressed visual information with the text modality, eliminating the need for the specialized text-vision aware cross-modal alignment module design[1, 12].
3.3 Temporal Modeling
When giving a sequence of video frames and a corresponding text input, VoCo-LLaMA, as introduced in Section 3.2, could serve as a vision encoding compressor to encode each frame. As shown in Figure 3, we obtain the compressed representation for each frame with the video frame tokens and the compression token . This yields a sequence of compressed video tokens, denoted by .
By harnessing the current capabilities of VoCo-LLaMA to interpret tokens derived from compressing individual frames, we can further enhance its functionality to handle temporal inputs. Specifically, we adopt a continue training process on VoCo-LLaMA by utilizing the time-series compressed token sequences , enabling it to capture and comprehend the temporal correlations among the compressed video tokens. With the inclusion of text tokens , VoCo-LLaMA learns the distribution . This training strategy builds upon existing image compression capabilities and allows VoCo-LLaMA to focus on temporal modeling, thereby streamlining the video understanding process.
3.4 Implementation Details
Regarding the training strategy and data, as mentioned earlier in Section 3.2, VoCo-LLaMA only requires learning to insert and compress VoCo tokens during the vision instruction tuning stage. We follow the common VLMs [2, 7] to encode the image input into vision tokens with vision encoder and a linear projector. We employ the pre-trained CLIP-ViT-L [40] as our visual encoder. For pre-trained large language models, we utilize Vicuna-7B [11]. Without introducing VoCo tokens, we first align the visual encoder and language model using the LLaVA-filtered CC3M [41] dataset with visual encoder and language model keeping frozen. During the instruction tuning phase of VoCo-LLaMA, incorporating multiple image understanding tasks is crucial for learning a scalable image compression model. Therefore, we construct the instruction pairs inspired by [12] using [7]. For video tuning, we further utilize WebVid [42] and the QA-pairs of Video-ChatGPT [39].
We conducted experiments on several common compression strategies with the same training setting and data for comparison. For the compression strategy with Q-Former, we employ the architecture in [1] and configure the query number to one, resulting in a single compression token. For the compression strategy with average pooling, we follow the design of the single content token in [12], as illustrated in Figure 1 (b). For more details on the implementation, please refer to the Appendix A.
4 Experiments
4.1 Datasets
In this work, we conduct experiments on several common visual understanding benchmarks for vision compression. In particular, we report results on GQA [43], MMB (MMBench) [44], MME [45], POPE [46], SEED-Bench [47], SQAI (Image-based setting in ScienceQA) [48] and VQAv2 (VQA V2) [49]. By observing the model’s performance on these image understanding benchmarks before and after compression (i.e. with initial vision tokens / VoCo tokens), we can observe the effects of the visual information loss that occurs during the vision compression process. We evaluate the performance on these visual understanding benchmarks in accordance with the details outlined in [2]. As for the video domain, we evaluate the zero-shot performance on several video question-answering benchmarks. MSVD-QA [50] is a video QA dataset consisting of 1,970 video clips with 50,505 QA pairs, built upon the Microsoft Research Video Description Corpus [51]. MSRVTT-QA [50] is a large-scale video QA dataset featuring 10K videos and 243K QA pairs with complex scenes, based on the MSR-VTT dataset [52]. ActivityNet-QA [53] is a fully annotated video QA dataset containing 58K QA pairs derived from 5,800 complex web videos from the ActivityNet dataset [54].
4.2 Vision Compression Configuration
In the primary experiment of vision compression, we present the results of compressing all vision tokens of an image into a single VoCo token. To rigorously quantify the performance loss of VoCo-LLaMA during compression, we designed two comparative training settings: the Upper Bound model, which represents the best compression performance. The ideal case for a visual compression model is to obtain the same visual understanding capability as the upper bound model. And the Lower Bound model, which represents the worst compression performance.
The initialization model is trained by integrating VoCo tokens in a manner analogous to VoCo-LLaMA, without modifying the attention mask strategy. During inference, we employ a standard causal attention mask. This setting effectively controls for performance fluctuations induced by the introduction of additional special tokens. In contrast, the random compression model is trained under identical settings as the initialization model. During inference, we restrict the visibility of text tokens to only the VoCo token, isolating the visual information. This setup represents a scenario without vision compression training, providing a baseline for evaluating. Based on the performance boundary model, the compression retention rate can be subsequently calculated as .
4.3 Results
Vision Compression. Table 1 presents the results of VoCo-LLaMA on vision compression. To explore the maximum potential of our method, we report the highest achievable compression ratio, which compresses vision tokens into a single VoCo token. We report results of our compression model on various common visual understanding benchmarks, as well as the compression retention rates defined based on upper and lower bound models introduced in Section 4.2. It can be observed that our method preserves the original visual information to a large extent, even at an extremely high compression ratio of 576. Specifically, we achieved an average compression retention rate of 83.7% across seven widely used benchmarks. Especially on MMBench, POPE and VQAv2, our method retained more than 85% of the performance during compression. The results indicate that VoCo-LLaMA can effectively compress vision tokens. Moreover, our method consistently outperforms the performance lower bound model of random compression across all benchmarks. This demonstrates that the advantages of VoCo-LLaMA, such as significant reductions in context length and improved calculation efficiency, outweigh any potential performance loss caused by compression.
We additionally compare our method with previous common learning-based approaches (i.e., Q-Former and average pooling) for vision token compression. Our method significantly outperforms previous methods across all benchmarks. Specifically, we observe an improvement of 19.6% in average compression retention rate, surpassing the average pooling compression strategy. In contrast, while Q-Former has demonstrated impressive capabilities in capturing visual features with 32 queries, its performance undergoes a substantial decline when the query count is reduced to a single digit. This proves that the compression method of VoCo-LLaMA, which utilizes the knowledge distillation from large language models itself, maintains more valuable vision information than that of average pooling or query-based compression.
| Method | Token | GQA | MMB | MME | POPE | SEED | SQAI | VQAv2 | Avg. |
| Upper Bound | 576 | 61.1 | 64.0 | 1487.2 | 85.0 | 57.9 | 66.5 | 77.7 | - |
| 100% | 100% | 100% | 100% | 100% | 100% | 100% | 100% | ||
| VoCo-LLaMA | 1 | 57.0 | 58.8 | 1323.3 | 81.4 | 53.7 | 65.4 | 72.3 | - |
| 82.5% | 87.5% | 81.2% | 88.4% | 80.0% | 81.0% | 85.2% | 83.7% | ||
| Avg. Pool [12] | 1 | 52.9 | 55.5 | 1210.3 | 79.1 | 50.3 | 62.2 | 65.0 | - |
| Linear | 65.0% | 79.6% | 68.1% | 81.0% | 63.8% | 25.8% | 65.2% | 64.1% | |
| Q-Former [1] | 1 | 51.1 | 51.7 | 1079.7 | 77.3 | 47.2 | 62.7 | 63.4 | - |
| 57.3% | 70.5% | 53.2% | 75.2% | 49.0% | 34.5% | 60.8% | 57.2% | ||
| Lower Bound | 1 | 37.7 | 22.3 | 617.3 | 53.9 | 36.9 | 60.7 | 41.2 | - |
| 0% | 0% | 0% | 0% | 0% | 0% | 0% | 0% |
| Token | MMB | GQA | VQAv2 | SEED | Avg. |
| 576 | 64.0 | 61.1 | 77.7 | 57.9 | 100% |
| 128 | 61.0 | 59.8 | 76.9 | 59.1 | 97.7% |
| 64 | 60.5 | 60.4 | 75.4 | 56.3 | 93.7% |
| 32 | 59.4 | 60.2 | 75.3 | 56.2 | 92.6% |
| 16 | 58.6 | 59.4 | 75.4 | 56.2 | 91.3% |
| 8 | 58.7 | 59.2 | 75.3 | 56.3 | 91.3% |
| 4 | 60.4 | 58.4 | 74.5 | 56.0 | 90.4% |
| 2 | 60.1 | 57.7 | 73.5 | 55.0 | 87.8% |
| 1 | 58.8 | 57.0 | 72.3 | 53.7 | 83.8% |
| 1 | 22.3 | 37.7 | 41.2 | 36.9 | 0% |
Number of VoCo tokens. We evaluate the impact of the number of VoCo tokens on vision compression performance. As shown in Table 2, we report the results of VoCo-LLaMA on four widely used benchmarks as the number of VoCo tokens increases from 1 to 128, along with the average compression retention rates across four benchmarks. Figure 4 illustrates the trend of compression performance retention as the number of VoCo tokens varies, where the green and red lines represent the upper and lower bounds of compression performance, respectively. We adopted the same training settings and data as in the main experiments.
It can be observed that as the number of VoCo tokens grows, the overall compression performance of the model shows an upward trend. Increasing the number of tokens within the range of fewer than 10 tokens results in a significant improvement in compression performance. Finally, when conducting 128 VoCo tokens, the model achieves an average compression performance retention rate of 97.7%, indicating that the performance loss due to compression is almost negligible when compressing into more than 100 tokens. Interestingly, we observe that when training with 128 VoCo tokens, the result on the SEED-Bench exceeds the initial performance upper bound of the model.
| Method | context | content | GQA | POPE | SQAI | VQAT |
| LLaMA-VID [12] | 1 | 16 | 58.2 | 83.1 | 67.4 | 50.8 |
| 1 | 4 | 56.2 | 83.5 | 68.7 | 49.1 | |
| 1 | 1 | 55.5 | 83.1 | 68.8 | 49.0 | |
| VoCo-LLaMA | 0 | 1 | 58.3 | 85.0 | 69.5 | 52.7 |
| VoCo-LLaMA + context [12] | 1 | 1 | 58.5 | 85.6 | 69.2 | 52.7 |
Method of Compression. We compare our method with LLaMA-VID on vision compression, specifically evaluating its full model that utilizes both context and content tokens. For a fair comparison, VoCo-LLaMA is trained under the exact same settings and applied the same visual encoder, EVA-G [55], in this experiment. As shown in Table 3, our method outperforms the previous approach across the board when using a single content compression token, even surpassing the performance of LLaMA-VID when it uses multiple context tokens. In particular, we could observe an improvement of 2.8 and 3.7 on GQA and VQAT benchmarks, respectively. Additionally, we examine the role of context tokens, which facilitate cross-modal interactions between vision tokens and text, in VoCo-LLaMA. The impact of context tokens is not as significant as demonstrated in LLaMA-VID, possibly as VoCo-LLaMA’s inherent design already facilitates cross-modal alignment.
Inference Efficiency. VoCo-LLaMA can bring significant efficiency and storage gains during the inference phase. Due to our model’s design, the representation of compressed image (i.e., transformer activations on top of VoCo tokens) can be stored and repeatedly utilized in the form of a KV cache. We conduct a comparative analysis of CUDA time, FLOPs, and KV Cache storage size during the inference process, and compare our method with the baseline method and the full caching method. The baseline method, as its name suggests, does not employ any caching strategy and directly encodes and infers images. In contrast, the full caching method stores the uncompressed transformer activations upon all vision tokens as KV caches. More specifically, we follow the approach of [56], storing the keys and values of each transformer layer. As displayed in Table 4, we conduct an inference efficiency analysis of three methods on a single NVIDIA A100-SXM4-40GB using identical lengths of text prompts and single-image inputs. Compared to the baseline model without caching strategy, VoCo-LLaMA achieves a significant reduction of 69.6% in CUDA time and 94.8% in FLOPs by caching compressed tokens with a KV Cache of length one. Relative to the full caching strategy, our method save 99.8% of cache storage while achieving lower CUDA time and FLOPs, demonstrating the inference efficiency gains brought by our vision compression approach. Please refer to Appendix B for further discussion and details for inference efficiency.
| Method | Token | KV Cache | Storage | CUDA | FLOPs | |||
| Length | Memory (MB) | Time (ms) | (T) | |||||
| Baseline | 576 | - | - | - | 440.5 | - | 9.6 | - |
| Full Caching | 576 | 576 | 302.4 | - | 154.9 | 64.8% | 1.2 | 87.5% |
| VoCo-LLaMA | 1 | 1 | 0.525 | 99.8% | 134.0 | 69.6% | 0.5 | 94.8% |
| Method | Visual | Res. | Image | MSVD-QA | MSRVTT-QA | ANet-QA | |||
| Encoder | Token | Acc | Score | Acc | Score | Acc | Score | ||
| Methods w/o Vision Compression | |||||||||
| FrozenBiLM [57] | CLIP-L | 224 | 256 | 32.3 | - | 16.8 | - | 24.7 | - |
| Video-LLaMA [35] | EVA-G | 224 | 256 | 51.6 | 2.5 | 29.6 | 1.8 | 12.4 | 1.1 |
| LLaMA-Adapter [58] | - | 224 | - | 54.9 | 3.1 | 43.8 | 2.7 | 34.2 | 2.7 |
| VideoChat [36] | - | 224 | - | 56.3 | 2.8 | 45.0 | 2.5 | 26.5 | 2.2 |
| Video-ChatGPT [39] | CLIP-L | 224 | 256 | 64.9 | 3.3 | 49.3 | 2.8 | 35.2 | 2.7 |
| BT-ADapter [37] | CLIP-L | - | - | 67.5 | 3.7 | 57.0 | 3.2 | 45.7 | 3.2 |
| Vista-LLaMA [59] | EVA-G | 224 | 256 | 65.3 | 3.6 | 60.5 | 3.3 | 48.3 | 3.3 |
| Chat-UniVi [34] | CLIP-L | 224 | 256 | 69.3 | 3.7 | 55.0 | 3.1 | 46.1 | 3.3 |
| Methods w/ Vision Compression | |||||||||
| LLaMA-VID [12] | EVA-G | 224 | 2 | 69.7 | 3.7 | 57.7 | 3.2 | 47.4 | 3.3 |
| VoCo-LLaMA | CLIP-L | 224 | 2 | 72.3 | 3.9 | 61.1 | 3.5 | 47.9 | 3.4 |
| 336 | 2 | 72.6 | 3.9 | 61.2 | 3.5 | 47.9 | 3.4 | ||
| 224 | 8 | 73.4 | 3.9 | 62.0 | 3.5 | 48.5 | 3.4 | ||
| 336 | 8 | 73.5 | 3.9 | 62.3 | 3.5 | 48.6 | 3.4 | ||
Video Understanding. We further evaluate the performance of VoCo-LLaMA on three widely used video understanding benchmarks, reporting results for input image resolutions of 224 and 336, respectively. First, we discuss the video understanding methods that utilize vision compression. Ensuring fair comparison, we adopted the same compression ratio as previous method [12], compressing each video frame into 2 VoCo tokens for training and inference in the video understanding model. Our method consistently outperforms previous video compression methods across all three benchmarks. Specifically, on the MSVD-QA and MSRVTT-QA datasets, VoCo-LLaMA achieved accuracies of 72.3% and 61.1%, respectively, corresponding to absolute gains of 3.7% and 5.9% over the previous best methods. Moreover, our method achieves the highest scores of 3.9 and 3.5, respectively.
In comparison to video understanding methods that do not employ vision compression, our approach, which represents each video frame with a mere 2 VoCo tokens, demonstrates strong competitiveness against methods that utilize 256 or more vision tokens per frame. To further explore the potential of VoCo-LLaMA, we opted to compress video frames into the number of VoCo tokens that exhibited the optimal compression performance within the 0 order of magnitude (i.e., 8 tokens). Notably, as we increase the number of tokens, our method effectively leverages additional visual information, resulting in improved performance. We also analyse the performance loss caused by the use of vision compression in video domain, as detailed in the Appendix D. In summary, VoCo-LLaMA capitalizes on the strengths of its efficient vision compression to achieve robust performance on video data, all while bypassing the intricate temporal modeling designs.
5 Limitations
VoCo-LLaMA presents a novel solution for vision compression, but several limitations remain. Our training paradigm, while effective in enabling the model to comprehend compression tokens, inadvertently diminishes its ability to understand uncompressed vision tokens. Moreover, our model struggles to effectively handle tokens with diverse fine-grained compression levels. Addressing these limitations could unlock the door to infinite-length input visual-language models.
6 Conclusion
In this paper, we propose VoCo-LLaMA, the first approach to compress visual information using large language models. By distilling the LLM’s understanding of vision tokens into a compact representation, we achieve efficient compression of vision tokens. Our method can compress hundreds of vision tokens into a single VoCo token, while minimizing visual information loss. VoCo-LLaMA significantly reduces cache storage and boosts efficiency during the inference stage. Moreover, our method exhibits exceptional performance in learning temporal understanding on video data with continuous training, showcasing its promising potential. In summary, our approach offers a promising solution for fully utilise the limited and valuable context window of VLMs, making them more scalable and efficient for various multi-modal applications.
Appendix A Additional Implement Details
A.1 Experiment Details
We elaborate on the training details and hyperparameter design of VoCo-LLaMA during the training process. Since our method only requires introducing VoCo tokens during the visual instruction tuning stage, the hyperparameters reported are specific to the fine-tuning stage. We followed [2, 7] to configure the majority of the model parameters. Table 6 provides a detailed list of hyperparameters involved in the model training phase. Under the aforementioned parameter settings, VoCo-LLaMA is fine-tuned on 8 NVIDIA A100-SXM4-40GB GPUs.
| Hyperparameter | Value |
| Overall batch size | 128 |
| Learning rate | 2e-5 |
| LR Scheduler | Cosine decay |
| DeepSpeed ZeRO Stage | ZeRO-3 |
| Optimizer | AdamW |
| Warmup ratio | 0.03 |
| Epoch | 2 |
| Weight decay | 0 |
| Precision | bf16 |
A.2 Instruction Tuning Details
In this subsection, we delve into the design of instructions and special tokens for VoCo-LLaMA during the instruction tuning stage, illustrating the details with concrete examples. As shown in LABEL:lst:app_list_1, by leveraging the pre-trained language model Vicuna[11] as the foundation for our method, we ensured a consistent system prompt and dialogue design to prevent the model from learning or adapting to new instruction formats during fine-tuning.
Furthermore, the VoCo-LLaMA architecture necessitates organizing the input sequence in a specific order, i.e., vision tokens, VoCo tokens, and text tokens. However, demonstrated in LABEL:lst:app_list_1 [Common VLM], the standard "Conversation" and "Input" formats reveal that the special token <image> and its corresponding encoded <Vision Tokens> are situated in the middle of the dialogue sequence. To address this, we made modifications to align them in [VoCo-LLaMA]. We relocated the special token <image> to the forefront of the input prompt and appended the <VoCo Tokens> for compression. Concurrently, we refined the system prompt to accommodate the adjusted vision token position, thereby enabling the LLM to better comprehend and model the location of vision token.
Appendix B Inference Efficiency Details
B.1 CUDA Time
We present a comparative efficiency analysis of VoCo-LLaMA, the baseline and full-caching strategies during the inference phase in Table 4. Here, we provide additional details on the CUDA time measurement during the inference phase. We primarily consider the following components that contribute to the reported CUDA time: image encoding time (if applicable), kv cache load time (if applicable), and transformers forward time. We exclude other computational times that are not dependent on the model itself and the caching strategy, such as model loading time, from the CUDA time measurement. For the baseline model, which is a regular visual language model, the inference process requires additional visual information encoding, and, due to the lack of caching, the model must compute the transformer activations and generate text on all tokens (uncompressed visual and textual tokens) during inference.
Compared to the Full Caching strategy, the improvement in CUDA time achieved by VoCo-LLaMA can be attributed to two primary factors: Firstly, the time required to load the kv cache is significantly reduced, owing to the smaller storage size and shorter read time resulting from the 576 compression multiplier achieved by our approach. Secondly, the transformers forward time is also improved, although this is largely dependent on the length of the new input token (i.e., text token), which is identical in both cases. However, VoCo-LLaMA’s shorter prefix enables more efficient computation, leading to faster processing times. Notably, since both approaches cache transformer activations on top of vision tokens or compressed tokens, the image encoding duration is not a contributing factor to the overall CUDA time.
B.2 Storage Memory and FLOPs
The storage memory reported in Table 4 corresponds to the full transformer activations on top of the vision tokens or compressed VoCo tokens. This also refers to the size of the kv cache loaded during the inference process. Notably, the storage precision is bf16, which is consistent with the precision used in our training process.
Appendix C Implementation Code
LABEL:lst:app_list_2 presents the concrete implementation of introducing VoCo tokens and the attention mask mechanism in VoCo-LLaMA, corresponding to Equations 2 and 3. To facilitate reader understanding, we provide a brief code snippet implemented using PyTorch, accompanied by annotations.
Appendix D Video Compression Analysis
We have analyzed the vision compression performance in Table 1, and in this subsection, we further analyze and discuss the video compression performance. The setup of the boundary model is similar to the one described in Section 4.2, with the exception that the vision token input from a single image is replaced by a sequence of vision tokens from multiple video frames. Considering methods such as [34], our final setup samples 32 frames of a video during both training and inference phases. To ensure that the context window of the boundary model can accommodate the input length of the video frame sequences, we employ a 224224 resolution input for the frames and enlarged the context window of its language model by a factor of four using [60].
As shown in Table 7, VoCo-LLaMA achieves nearly lossless compression of video frames at a rate of 256, achieving an average compression multiplier of 97.4% on three widely-used video QA benchmarks. As the input image count increases, the vision compression operation of VoCo-LLaMA can reduce the total length of the vision token, mitigate the loss of information at a distance, and bring further gains to visual understanding.
| Method | Token | MSVD-QA | MSRVTT-QA | ANet-QA | Avg. | |||
| Acc | Score | Acc | Score | Acc | Score | |||
| Upper Bound | 256 | 68.8 | 3.7 | 57.3 | 3.2 | 46.5 | 3.3 | - |
| 100% | 100% | 100% | 100% | 100% | 100% | 100% | ||
| VoCo-LLaMA | 1 | 68.1 | 3.7 | 56.6 | 3.1 | 45.9 | 3.3 | - |
| 97.2% | 100% | 97.5% | 92.3% | 97.1% | 100% | 97.4% | ||
| Lower Bound | 1 | 43.5 | 2.2 | 29.4 | 1.9 | 25.9 | 2.2 | - |
| 0% | 0% | 0% | 0% | 0% | 0% | 0% | ||
Appendix E Visualized Examples of Output
Figure 5 illustrates the visualized examples of VoCo-LLaMA for vision compression. We compare our approach, which employs a single VoCo token, with the Upper Bound model described in Section 4.2 (denoted as U.B. throughout the figure). Correct responses are highlighted in green, while incorrect responses are marked in red. We further show some of the failure cases. For questions that require attention to finer details of the input image, such as the third question in the second case, VoCo-LLaMA sometimes produces incorrect descriptions. Notably, when the VoCo token count increases to 8, these erroneous outputs are significantly improved, as the compression retention rate rises.
Appendix F Broader Impacts
Our research relates to the Large Language Models and Large Vision Language Models. Although the training data we employ is public and compliant, there are potential negative societal impacts to consider. One possible concern is the generation of unintended or malicious outputs, which could be used to spread disinformation. To mitigate these risks, we propose that future work focus on developing mechanisms for monitoring and detecting misuse, as well as implementing gated release strategies for models to ensure they are used responsibly.